# Dual Alignment Maximin Optimization for Offline Model-based RL

### v2

@@ 5 Dual Alignment Maximin Optimization
Inspired by the analysis above, we propose Dual Alignment Maximin Optimization (DAMO), a unified approach to mitigate both OOD actions and states. We further explore the distinct roles of the inner minimization and the outer maximization in addressing the distribution shift challenge.
5.1 Maximin Objective for Unified Shift Mitigation
@@
We start from the root cause of distribution shifts, i.e., the discrepancy between learned and behavior policy behaviors in the underlying environment. To tackle this, we introduce a regularized objective that constrains the difference between these behaviors while optimizing for expected rewards:
@@
where $\alpha$ is a regularization hyperparameter. Next, we self-consistently incorporate synthetic data through a divergence upper bound, deriving the following surrogate objective:
@@
This surrogate objective integrates two essential elements: a data alignment regularizer $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to ensure compatibility between synthetic and offline data, and a behavior alignment regularizer $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ that maintains consistent policy performance across models and real environments.
While this objective (1) theoretically addresses the synthetic-to-real distribution mismatch (as shown later in Sec. 5.3), it is impractical to directly compute due to the inaccessibility of $\rho_{T}^{\pi}$. To make this objective tractable, we apply the convex conjugate theorem into the dual formulation of reinforcement learning, reformulating it as a maximin optimization problem. This reformulation forms the core of our proposed DAMO, a unified maximin framework for mitigating distribution shifts in offline model-based RL.
Theorem 5.1.
@@
$\max\limits_{\pi}\min\limits_{\tilde{Q}(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)],$ (2)
where $f_{\star}$ is the conjugate function of the convex function $f$, and $\Phi(s,a,s^{\prime})=\log r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)$, with $\tau^{\pi}Q(s,a)=\gamma\sum_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})$. Here, $\mu_{0}$ denotes the distribution of the initial state, and $\alpha$ is a hyperparameter to control the degree of conservatism.
5.2 Practical Implementation of DAMO
To implement the maximin optimization (2) practically, we employ a classifier to approximate the data alignment term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$. Specifically, we train a classifier $h(s,a,s^{\prime})$ using the following loss function to distinguish transitions sampled from offline data and those generated by the model:
@@ 5.2 Practical Implementation of DAMO
Here, $\mathcal{D}_{R}$ and $\mathcal{D}_{M}$ represent the offline and the model-generated data, respectively. Based on the learned classifier $h^{\star}(s,a,s^{\prime})$, the data alignment modification $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ can be computed using the following analytical expression:
@@
Note that (2) can be divided into two phases: an inner value estimation phase and an outer policy improvement phase, resembling the structure of the actor-critic framework. This structure is directly adapted to implement DAMO. The complete DAMO framework is outlined in Algorithm 1.
Algorithm 1 DAMO
0:  Offline dataset $\mathcal{D}_{R}$, dynamics model $M$, actor network $\pi_{\phi}$, critic network $Q_{\theta}$, classifier $h_{\psi}$.
1:  Train the probabilistic dynamics model $M$ on $\mathcal{D}_{R}$
2:  Initialize the model dataset buffer $\mathcal{D}_{M}\leftarrow\varnothing$
3:  for each epoch do
4:     Generate synthetic rollouts by model $M$
5:     Add transition data in these rollouts to $\mathcal{D}_{M}$
6:     Update classifier $h_{\psi}$ according to Eq. (3)
7:     Calculate $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ according to Eq. (4)
8:     Update critic $Q_{\theta}$ via minimizing Eq. (2)
9:     Update actor $\pi_{\phi}$ via maximizing Eq. (2)
10:  end for
5.3 Theoretical Insights for DAMO
The core structure of DAMO involves mapping the inner minimization in (2) to the estimation of a dual conservative value $\tilde{Q}(s,a)$, and the outer maximization to policy improvement based on this value. In this subsection, we present deeper theoretical insights into the workings of DAMO.
Dual Conservative Value Estimation. We first demonstrate the existence of both explicit and implicit reward penalties in the optimization objective (2), which contribute to the construction of the dual conservative value estimation.
Theorem 5.2.
The optimal solution $Q_{\pi}^{\star}$ for the inner minimization optimization in (2) satisfies the following equation:
$Q_{\pi}^{\star}=r-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q_{\pi}^{\star}.$
This indicates that DAMO inherently introduces an explicit data alignment modification, $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$, and an implicit behavior alignment adjustment, $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$. Minimizing the inner objective effectively reshapes the reward function, $r(s,a,s^{\prime})$, into a refined reward, $\tilde{r}(s,a,s^{\prime})$, which represents the dual conservative value estimation:
@@ 5.3 Theoretical Insights for DAMO
The data alignment modification $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$ penalizes transition $(s,a,s^{\prime})$ that exhibit discrepancies between synthetic and offline data, promoting alignment with in-distribution transitions. This helps ensure accurate value estimation for these transitions. Meanwhile, the behavior alignment adjustment $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$ focuses on that different dynamics impact the distribution of $(s,a,s^{\prime})$. By penalizing these discrepancies, we encourage the agent to select policies that exhibit similar performance in both the underlying environment and the dynamics model, thus mitigating the impact of OOD states and ensuring consistent policy behavior between dynamics models and real environments.
Consistent Policy Improvement. Unlike conventional offline model-based approaches that optimize policies purely by maximizing estimated values, DAMO ensures the consistency between value estimation and policy improvement. Specially, DAMO employs $\tilde{Q}(s,a)$ to minimize the objective (2) and utilizes $\pi$ to maximize it. Conventional approaches typically follow a bilevel optimization framework, while DAMO employs a maximin paradigm, ensuring consistency and preserving dual conservatism. As demonstrated in Sec. 6.1, this consistent structure maintains the conservatism of value estimation during policy improvement.
Once the inner minimization in (2) is solved, the outer maximization problem is equivalent to solving the surrogate objective (1), which provides a lower bound for the standard RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$:
Theorem 5.3.
(1) is a lower bound of $\mathcal{J}(\pi)$ for all $\pi$:
@@
This demonstrates that the equivalence between objective (2) and the standard reinforcement learning objective ensures that the learned policy performs well in the environment.
6 Experiments
In this section, we focus on the following key aspects: 1) The capability of DAMO to handle synthetic-to-real distribution mismatch, as well as the contributions of its components. 2) The comparative performance of DAMO against existing methods on standard offline RL benchmarks. 3) The significant impact of hyperparameter settings and implementation approaches on the overall performance of DAMO.
We delve into these aspects using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012). Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of DAMO are consistent with the settings of this library. The details of each experiment and the specific hyperparameters configuration can be found in Appendix D.
6.1 Effectiveness of DAMO Paradigm
In this section, we present empirical evidence demonstrating the dual capabilities of DAMO: 1) effectively aligning model-generated data with offline data distributions, and 2) consistently maintaining policy stability between dynamics models and real environments. Our detailed ablation study thoroughly investigates two key components of DAMO: inner value estimation and outer policy improvement.
@@ 6.1 Effectiveness of DAMO Paradigm
Figure 3: (a) (b) (c) illustrate the results of DAMO without the data alignment term (w/o er), DAMO without the behavior alignment term (w/o ir), DAMO. The left panels depict the distributions of state-action pair $(s,a)$ for offline and synthetic data. The right panels display the distributions of states for real and synthetic data.
(a) Normalized Score
(b) Estimated Value
Figure 4: Training process of DAMO, w/o ir, w/o er. (a) The data alignment term is the primary factor that influences policy improvement, while the absence of the behavior alignment term results in performance degradation in the final stages. (b) The lack of the data alignment term leads to the overestimation of the value.
6.1.1 Dual Conservative Value Estimation
In this part, we present an ablation study examining the individual contributions of the data alignment and behavior alignment terms. We systematically evaluate their impact by removing each term from DAMO and training policies on the hopper-medium-expert-v2 benchmark. To provide intuitive insights, we first visualize the state-action distributions of the final trained policies across three distinct dataset categories. These categories are defined as follows:
  • Offline Data: The transition pairs $(s,a,s^{\prime})$ are collected with behavior policy $\pi_{\beta}$ in dataset Hopper-medium-expert.
  • Synthetic Data: The transition pairs are generated by conducting training policy $\pi$ within the dynamics model $M$.
  • Real Data: The transition pairs are generated by conducting training policy $\pi$ within the real environment $T$.
Effectiveness of Data Alignment. To evaluate synthetic and offline data compatibility during training, we analyze the state-action pair distributions for both synthetic data and offline data under three distinct conditions: the complete DAMO framework, the configuration without the behavior alignment term (w/o ir), and the configuration without the data alignment term (w/o er). The visualization results are presented in the left column of Fig. 3. Building on these initial observations, we then examine the training dynamics of DAMO under three specific conditions to further investigate their significant impacts. The results are presented in Fig. 4. A comparative analysis between Fig. 3(a) and Fig. 3(c) reveals that the w/o er configuration fails to establish proper alignment between synthetic data and offline data, thereby inadequately mitigating OOD state-action pairs. The results in Fig. 3(b) further highlight that the data alignment term enables w/o ir configuration to align synthetic data with offline data. To clarify the detrimental effects of incompatibility between these two data sources, Fig. 4(b) illustrates that w/o er overestimates the value of OOD actions, leading to suboptimal training performance. This is corroborated by the poor evaluation scores of w/o er depicted in Fig. 4(a).
Effectiveness of Behavior Alignment. To assess the model-environment consistency during the testing phase, we follow a similar evaluation procedure. We visualize the state distributions of real environment and model-generated data under the same conditions. The visualization results are presented in the right column of Fig. 3. A comparative analysis between Fig. 3(b) and Fig. 3(c) reveals that while w/o ir successfully achieves synthetic-offline data compatibility and mitigates OOD actions, the absence of the behavior alignment term leads to incomplete mitigation of OOD states, manifesting policy inconsistency between dynamics model and real environment. Although the behavior alignment term imposes constraints, the results in Fig. 3(a) demonstrate that w/o er fails to effectively mitigate OOD states. The observed discrepancies in state distributions between real environment data and model-generated data under the w/o er highlight the critical importance of establishing robust compatibility between synthetic and offline data. This compatibility is essential for ensuring consistent policy performance across dynamics models and real environments. The influence is reflected in Fig.4(a), which shows performance degradation of w/o ir in the final stage.
6.1.2 Towards Consistent Policy Improvement
Table 1: The average value of OOD state-action pairs, $\infty$ for severe overestimation. The smallest values are bolded, which indicates the best mitigation of the overestimation problem.
halfcheetah hopper walker2d
Consistent Version 95.2 213.0 231.7
Inconsistent Version 101.4 219.4 $\infty$
$Real\ value$ 819.6 269.6 299.7
This section presents a statistical analysis of the value estimation performance of DAMO across multiple datasets, with comparisons against the true value function. Our results validate that the compatible policy improvement framework effectively maintains the conservatism of value estimation.
To investigate the significance of objective alignment in DAMO, we conduct a controlled comparison experiment by directly integrating DAMO with standard off-policy mechanisms (Inconsistent Version) while evaluating the critic’s estimation of state-action pair values and comparing it with the original DAMO (Consistent Version). We perform this analysis on the medium-expert dataset across three distinct environments: HalfCheetah, Hopper, and Walker2d. The obtained quantitative results are summarized in Table 1. The experimental findings reveal that the Consistent Version achieves conservative value estimations across all environments. Although the Inconsistent version maintains conservatism in the HalfCheetah and Hopper environments, it exhibits significant overestimation in the Walker2d environment. These results collectively demonstrate the critical role of objective alignment in consistently ensuring conservative value estimation and preventing overestimation errors.
@@ 6.1.2 Towards Consistent Policy Improvement
Table 2: Results on the D4RL Gym benchmark. The numbers reported for DAMO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. Since we haven’t obtained the results of O-DICE on random datasets, we didn’t include it in the table and used a dash (-) to indicate the absence of data. The top three scores are highlighted in bold. We use an asterisk (*) for the best score, an underscore (_) for the second, and bold text for the third.
@@
Task Name Algae Opti O-DICE MOPO COMBO TT RAMBO MOBILE DAMO (Ours)
@@
halfcheetah-medium-expert -0.8 91.1 93.2 83.6 90.0 95.0 95.4 101.4 103.1 $\pm$ 0.9
@@
6.2 Comparison Results

### v3

@@ 5. Dual Alignment Maximin Optimization
Inspired by the analysis above, we propose Dual Alignment Maximin Optimization (DAMO), a unified maximin approach to mitigate both OOD actions and states through dual alignment. We further explore the distinct roles of the inner minimization and the outer maximization of DAMO in addressing the distribution shift.
5.1. Maximin Objective for Shift Mitigation
@@
We start from the root cause of distribution shifts, i.e., the discrepancy between learned and behavior policy behaviors in the underlying environment $T$. To tackle this, we imitate the behavior regularization methods and introduce a regularized objective that constrains the difference between distributions of learned and behavior policy while optimizing for expected rewards:
@@
where $\alpha$ is a regularization hyperparameter. Next, we self-consistently incorporate synthetic data through establishing an upper bound of KL divergence, deriving the following surrogate objective:
@@
Where $f$ is a convex function that satisfies the inequality: $f(x)\geq x\log x,\forall x>0.$ This surrogate objective integrates two essential elements: a data alignment regularizer $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to ensure compatibility between synthetic and offline data, and a behavior alignment regularizer $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ that maintains consistent policy performance across dynamics models $M$ and real environments $T$.
While objective (1) theoretically addresses the synthetic-to-real distribution mismatch (as shown later in Sec. 5.3), it is impractical to compute due to the inaccessibility of distribution $\rho_{T}^{\pi}$. To address this problem, we apply the convex conjugate theorem to the dual formulation of the RL objective, reformulating it as a maximin optimization problem. This reformulation forms the core of DAMO, a unified maximin framework for mitigating the OOD issue.
Theorem. 1.
@@
$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)],$ (2)
where $f_{\star}(x)$ is the conjugate function of the convex function $f(x)$, and $\Phi(s,a,s^{\prime})=r-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\sum\limits_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})-Q(s,a)$
The proof of Theorem 1 utilizes the DICE to correct the discrepancies between distributions and can be found in Appendix A.2.
5.2. Practical Implementation of DAMO
To implement the maximin optimization (2) practically, we employ a classifier to approximate the data alignment term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$. We train a classifier $h(s,a,s^{\prime})$ with the following loss function to distinguish transitions sampled from offline data and synthetic data:
@@ 5.2. Practical Implementation of DAMO
$\mathcal{D}_{R}$ and $\mathcal{D}_{M}$ represent the buffer storing offline and synthetic data, respectively. Based on the learned classifier $h^{\star}(s,a,s^{\prime})$, $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ can be computed using the following analytical expression:
@@
Note that the maximin optimization problem (2) can be decomposed into two phases: the inner value estimation and the outer policy improvement, resembling the actor-critic framework. Thus, we adopt the actor-critic structure to facilitate the implementation of DAMO. The complete framework is listed in Appendix C.1.
5.3. Theoretical Insights for DAMO
The core structure of DAMO involves mapping the inner minimization in (2) to the estimation of a dual conservative value $Q(s,a)$, and the outer maximization to policy improvement based on the estimated conservative value. In this subsection, we present deeper theoretical insights into the working mechanism of DAMO.
Dual Conservative Value Estimation. We first show that the inner minimization within (2) includes both: (1) an explicit penalty that directly aligns the synthetic data with offline data, and (2) an implicit penalty that emerges during minimization, ensuring policy behavior consistency as the solution converges to the optimum.
Theorem. 2.
The optimal solution $Q_{\pi}^{\star}$ for the inner minimization optimization in (2) satisfies:
$Q_{\pi}^{\star}=r-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q_{\pi}^{\star}.$
The proof of Theorem 2 is provided in Appendix A.3. This theorem indicates that DAMO inherently introduces an explicit data alignment regularizer, $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$, and an implicit behavior alignment adjustment, $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$. The inner minimization thus reshapes the vanilla reward $r(s,a,s^{\prime})$ into a refined reward $\tilde{r}(s,a,s^{\prime})$:
@@ 5.3. Theoretical Insights for DAMO
The data alignment regularizer $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$ penalizes transition $(s,a,s^{\prime})$ that exhibit discrepancies between synthetic and offline data, promoting alignment with in-distribution transitions. This alignment helps ensure accurate value estimation for these transitions. Meanwhile, the behavior alignment adjustment $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$ focuses on that different dynamics impact the distribution of $(s,a,s^{\prime})$. By penalizing these discrepancies, we encourage the agent to select policies that exhibit similar performance in both the dynamics model and the underlying environment, thus mitigating the impact of OOD states during deployment and ensuring consistent policy behavior between dynamics models and real environments.
Consistent Policy Improvement. Unlike conventional offline actor-critic frameworks that optimize policies solely by maximizing estimated values, DAMO ensures the objective consistency between value estimation and policy improvement. Specifically, DAMO employs the conservative Q-value to minimize the objective (2) and utilizes $\pi$ to maximize it. Prior actor-critic approaches typically follow a bilevel optimization framework, while DAMO employs a maximin paradigm, ensuring objective consistency. As demonstrated in Sec. 6.1, this objective-consistency maintains the conservatism of value estimation during policy improvement.
Once the inner minimization is solved, the outer maximization reduces to the surrogate objective (1), providing a lower bound for the standard RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$:
Theorem. 3.
The surrogate objective (1) serves as a lower bound of $\mathcal{J}(\pi)$ for all $\pi$:
@@
The detailed proof of Theorem 3 can be found in Appendix A.1. This theorem demonstrates the equivalence between maximin objective (2) and the standard RL objective, thereby offering a theoretical guarantee that the policy learned through objective (2) can achieve consistent performance in the real environment.
6. Experiments
In this section, we focus on the following aspects: 1) The capability of DAMO to handle distribution mismatch, and the contributions of each component of DAMO. 2) The performance of DAMO against existing methods on standard benchmarks. 3) The impact of hyperparameter settings and implementation approaches. 4) The OOD generalization ability of DAMO in the task with relabeled data.
We delve into these aspects using the D4RL Fu et al. (2020) and NeoRL Qin et al. (2022) benchmark on the MuJoCo simulator Todorov et al. (2012). Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit. The basic parameters of DAMO are consistent with the settings of the library. The experiment details and the hyperparameter configuration is listed in Appendix E.
(a) Normalized Score
(b) Estimated Value
Figure 3. Training process of DAMO, w/o ir, w/o er. (a) The data alignment term is the primary factor that influences policy improvement, while the absence of the behavior alignment term results in performance degradation in the final stages. (b) The lack of the data alignment term leads to the overestimation of the value.
6.1. Effectiveness of DAMO Paradigm
In this section, we present empirical evidence demonstrating the dual capabilities of DAMO: 1) effectively aligning synthetic data with offline data, and 2) consistently maintaining policy stability between the dynamics models and real environments. Our detailed ablation study thoroughly investigates two key components of DAMO: inner value estimation and outer policy improvement.
6.1.1. Dual Conservative Value Estimation
In this part, we present an ablation study examining the contributions of the data alignment regularizer and behavior alignment regularizer. We evaluate their impact by removing each term from DAMO and training policies on the hopper-medium-expert task. Specifically, we test DAMO under three distinct conditions: the configuration without the data alignment regularizer (w/o er), the configuration without the behavior alignment regularizer (w/o ir), and the complete DAMO framework. The empirical results are presented in Fig. 3. To further provide intuitive insights, we present the visualization of distributions under each setting in Fig 4.
@@ 6.1.1. Dual Conservative Value Estimation
Figure 4. (a)-(c) illustrate the results of DAMO without the data alignment term (w/o er), DAMO without the behavior alignment term (w/o ir), and standard DAMO. Left panel: state-action (s, a) distributions (offline vs. synthetic). Right panel: state distributions $s$ (real vs. synthetic).
@@
Table 1. Results on the D4RL Gym benchmark. The numbers reported for DAMO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. Since we haven’t obtained the results of O-DICE on random datasets, we didn’t include it in the table and used a dash (-) to indicate the absence of scores. The top three scores are highlighted in bold. We use an asterisk (*) for the best score, an underscore (_) for the second, and bold text for the third.
@@
Task Name Algae Opti O-DICE MOPO COMBO $T$ RAMBO MOBILE DAMO (Ours)
@@
halfcheetah-medium-expert -0.8 91.1 93.2 83.6 90.0 95.0 95.4 101.4 103.1 $\pm$ $\textbf{0.9}^{*}$
@@
Effectiveness of Data Alignment. A comparative analysis between Fig. 4(a) and Fig. 4(c) reveals that w/o er fails to establish alignment between synthetic and offline data, thereby inadequately mitigating OOD state-action pairs. The results in Fig. 4(b) further highlight that the data alignment regularizer enables the w/o ir configuration to align synthetic data with offline data. To clarify the detrimental effects of incompatibility between these two data sources, the result in Fig. 3(b) illustrates that the w/o er configuration overestimates the value of OOD actions, leading to suboptimal training performance. This is corroborated by the poor evaluation score of the w/o er configuration depicted in Fig. 3(a).
Effectiveness of Behavior Alignment.  A comparison between Fig. 4(b) and Fig. 4(c) reveals that while w/o ir successfully achieves synthetic-offline data compatibility and mitigates OOD actions, the absence of the behavior alignment term leads to incomplete mitigation of OOD states, manifesting policy inconsistency between dynamics models and real environments. Although the behavior alignment term imposes constraints, the results in Fig. 4(a) demonstrate that w/o er fails to effectively mitigate OOD states. The observed discrepancies in state distributions between real environment data and model-generated data under the w/o er highlight the critical importance of establishing robust compatibility between synthetic and offline data. This compatibility is essential for ensuring consistent policy performance across dynamics models and real environments. The influence is reflected in Fig.3(a), which shows performance degradation of w/o ir in the final stage of training.
6.1.2. Towards Consistent Policy Improvement
This section presents a statistical analysis of the value estimation performance of DAMO across multiple datasets, with comparisons against the true value function of the policy in the real environment. Our results validate that the compatible policy improvement framework effectively maintains the conservatism of value estimation.
To investigate the significance of objective alignment in DAMO, we conduct a comparison experiment by directly integrating DAMO with standard off-policy mechanisms (Inconsistent Version) while evaluating the critic’s estimation of state-action pair values and comparing it with the original DAMO (Consistent Version). The detailed implementation of the Inconsistent Version can be found in Appendix C.3. We perform this analysis on the medium-expert dataset across three distinct environments: HalfCheetah, Hopper, and Walker2d. The obtained quantitative results are summarized in Table 2. The experimental findings reveal that the Consistent Version achieves conservative value estimations across all environments. Although the Inconsistent version maintains conservatism in the HalfCheetah and Hopper environments, it exhibits significant overestimation in the Walker2d environment. These results demonstrate the critical role of objective alignment in persisting conservative value estimation and preventing overestimation.
Table 2. The average value of OOD state-action pairs, $\infty$ for severe overestimation. The smallest values are bolded, which indicates the best mitigation of the overestimation problem.
halfcheetah hopper walker2d
Consistent Version 95.2 213.0 231.7
Inconsistent Version 101.4 219.4 $\infty$
RealvalueReal\ value 819.6 269.6 299.7
6.2. Comparison Results
Table 3. Results on the NeoRL benchmark. We report the normalized scores averaged over the trained policy across 4 seeds. The results of MOBILE are obtained by executing the official implementation.
Task Name BC CQL TD3+BC EDAC MOPO MOBILE DAMO (Ours)
Halfcheetah-L 29.1 38.2 30.0 31.3 40.1 54.8 $47.55\pm 3.11$
Hopper-L 15.1 16.0 15.8 18.3 6.2 18.2 24.05 $\pm$ 2.18
Walker2d-L 28.5 44.7 43.0 40.2 11.6 23.7 $\textbf{46.59}\pm\textbf{9.18}$
Halfcheetah-M 49.0 54.6 52.3 54.9 62.3 79.0 $69.04\pm 1.13$
Hopper-M 51.3 64.5 70.3 44.9 1.0 43.4 $64.57\pm 9.17$
Walker2d-M 48.7 57.3 58.6 57.6 39.9 60.1 $\textbf{66.49}\pm\textbf{2.28}$
Halfcheetah-H 71.3 77.4 75.3 81.4 65.9 71.8 $67.23\pm 17.5$
Hopper-H 43.1 76.6 75.3 52.5 11.5 42.3 $74.87\pm 6.39$
Walker2d-H 72.6 75.3 69.6 75.5 18.0 71.9 $74.28\pm 2.03$
6.2.1. D4RL